# Position: AI Competitions Provide the Gold Standard for Empirical Rigor in GenAI Evaluation

### v1

@@ Abstract
In this position paper, we observe that empirical evaluation in Generative AI is at a crisis point since traditional ML evaluation and benchmarking strategies are insufficient to meet the needs of evaluating modern GenAI models and systems. There are many reasons for this, including the fact that these models typically have nearly unbounded input and output spaces, typically do not have a well defined ground truth target, and typically exhibit strong feedback loops and prediction dependence based on context of previous model outputs. On top of these critical issues, we argue that the problems of leakage and contamination are in fact the most important and difficult issues to address for GenAI evaluations. Interestingly, the field of AI Competitions has developed effective measures and practices to combat leakage for the purpose of counteracting cheating by bad actors within a competition setting. This makes AI Competitions an especially valuable (but under-utilized) resource. Now is time for the field to view AI Competitions as the gold standard for empirical rigor in GenAI evaluation, and to harness and harvest their results with according value.
@@ 1 Introduction
As Generative AI (GenAI) models such as Large Language Models (LLMs) become ever more important to the field and to the world at large, it has become clear that performing empirical evaluation of these models and methods is extremely difficult to do in a rigorous and comprehensive way. This difficulty is of course not due to lack of effort or expertise by researchers. Indeed, enormous effort and resources have been poured into creating myriad benchmarks and test cases (Chiang et al., 2024b; Fourrier et al., 2024; Hendrycks et al., 2021; Cobbe et al., 2021; Zellers et al., 2019; Chen et al., 2021b). However, even accounting for these many important efforts and achievements, our position is that the current state of evaluation is insufficient to meet the needs of this moment in GenAI for the field and for the world.
In our view, the root cause of this insufficiency is that the evaluation needs of GenAI models fundamentally break the paradigm of traditional benchmarking that served the field of machine learning (ML) so well during decades of remarkable progress. This breakage goes beyond the familiar difficulty of defining what, exactly, is in the training data for an LLM. In our view, we need a broader conception of generalization for GenAI that moves beyond the idea of generalizing to new independently drawn examples from a stationary distribution, and instead refers to performing well on tasks that are entirely novel from a model’s perspective. This higher bar is rooted in commonsense standards for human intelligence (Chollet, 2019; Dennett, 1991), but has far reaching consequences, most notably that it implies that the problems of data leakage and contamination in evaluation are the most pressing concerns.
Together, these factors imply that rigorous and robust evaluation GenAI models requires a steady source of novel tasks structured to avoids leakage, contamination, and other forms of inadvertent “cheating”. Fortunately, AI Competitions—such as those hosted on Kaggle and similar platforms—offer a ready-made solution, providing a continual source of new tasks for evaluation and significant structures to avoid leakage and related issues.
@@ 1.1 Summarizing Our Position
  • GenAI evaluations should be considered leaked the moment it has been shared online or sent over the wire.
@@
  • Meta-analyses should be valued as highly in the field of AI as they are in fields such as medicine.
@@ 1.2 Structure of this paper
In the remainder of this paper, we will first review the most typical structure and assumptions in traditional ML evaluation and discuss why it is insufficient for GenAI evaluations. We will examine the nature of generalization for GenAI, how this leads to specific concerns around leakage, and additionally show how goals of reproducibility and robustness in evaluation may be fundamentally at odds. We will then show how difficult the problem of leakage is even for traditional ML evaluations with some brief case studies, and look at current GenAI benchmarks that are aiming to overcome leakage and contamination. We finish with an examination of the ways that AI Competitions address these issues, discuss our recommendations, and examine alternate viewpoints. Our goal is to provide convincing support of the view that AI Competitions do indeed provide a gold standard for empirical rigor in evaluating GenAI models, and that the field should place accordingly high value and attention on their results.
Figure 1: IID Evaluations vs. Novelty-Centric Evaluations. In the IID evaluation, left, both training and test data are drawn from the same distribution, resulting in significant overlap in examples in each set. In the novelty-centric version, right, no test example is allowed to be too similar to any given training example. We argue that this conceptualization more closely mirrors desired behavior for GenAI evaluations, where generalization is expected to connote the ability to respond well on totally novel inputs.
@@ 2 Background: Revisiting Benchmarking
Traditional ML benchmarking has been founded on the idea of a test-train split, in which an evaluation is structured by training a model from scratch on a given portion of training data and then evaluating that trained model on a holdout set of test data (Mitchell, 1997). This conceptual structure is so fundamental to modern ML practice that it may sometimes be taken for granted and not carefully examined. So let us take a moment to reflect this basic structure and its implications.
In classical supervised ML, the most common traditional setup is to evaluate a model $f({\bf x})\rightarrow y$, with ${\bf x}\in\Re^{d}$ as feature vectors in some $d$ dimensional feature space and $y\in Y$ as a space of possible labels, such as $\{0,1\}$ for binary classification or $(0,1)$ for regression on probabilities. Labeled examples $({\bf x},y)$ are assumed to have come from some distribution $D$. The training set $D_{train}$ and test set $D_{test}$ are each independently and identically drawn (IID) from $D$, with only the examples in the training set used to fit the model $f({\bf x})$ and only the examples in the test set used to evaluate the model (Mitchell & Mitchell, 1997).
The IID requirement on test-train splits is often taken as a footnote in practice, but in reality it is a cornerstone of the robustness of this setup. The reason for this is that we fundamentally wish our evaluations to be interpretable as statements on generalization ability of our models: we wish to know how the model will perform on future, previously unseen data. But achieving this is harder than it may sound, because the ML models in question are often of extremely high dimensionality and thus may be prone to overfitting.
@@
A second approach is to use additional data for evaluation. The issue here to be aware of is the classical statistical trap that correlation does not necessarily imply causation, and that trying to assess generalization ability of a model that has no specific mechanism for disambiguating correlation from causal factors may lead to wildly unreliable performance estimates. It is this issue that the IID assumption addresses—when we know that all test data is drawn IID from the same distribution as the training data, we know that all correlations that held at training time will reappear with the same characteristics at test time, and thus we can take performance on holdout test data as a reasonable estimate of generalization ability. The IID assumption in many ways enabled modern machine learning to advance as a research field, because it forms the theoretical underpinning of all evaluations. And indeed, it is a truism that moving ML models from research to deployed production is difficult precisely because of the fact that the IID assumption often does not hold in practice (Chen et al., 2021a).
@@ 2.1 The Rise of Reproducible Benchmarks
One statistical shortcut that immediately became standard practice was instead of drawing a new training set and test set from $D$ for every new evaluation, researchers would make one draw of each and use them as a canonical train / test set. The primary benefit of this approach (besides convenience) was that these paired test-train splits could now be used as reproducible benchmarks. All future researchers could replicate the exact problem setup, giving a new untrained model the same training data for training and the same test data for evaluation, allowing for full apples-to-apples comparisons. This approach was wildly successful, with canonical benchmarks such as MNIST (LeCun & Cortes, 2010) and ImageNet (Deng et al., 2009) responsible for driving incredibly rapid progress in computer vision, for instance, and benchmarks such as (Rajpurkar et al., 2016; Marcus et al., 1993; Diemert Eustache, Betlei Artem et al., 2018) moving forward email spam classification, natural language processing, and many others. Websites such as the UCI repository (Kelly et al., ) and OpenML (Vanschoren et al., 2014) among many others have been remarkably valuable to the field for this reason.
@@ 2.2 Surprisingly, Overfitting Was Not the Main Issue
Given that one of the fundamental concerns in evaluating models was ensuring that we could properly address generalization and avoid overfitting, it is reasonable to ask whether having a field re-use the same standard benchmark datasets in thousands of papers might not lead to overfitting. As authors, we were deeply surprised by the work of (Roelofs et al., 2019b) showing that in practice, this appears to actually not have been a problem. In (Recht et al., 2019), these authors shows that the rank ordering of ImageNet models when evaluated on brand new data was remarkably consistent with the rank ordering of those models on the benchmark data, despite its wide reuse. And in follow-up work, it was shown that similarly evaluation on public leaderboard data on Kaggle competitions was a remarkably good indicator of rank ordering on private holdout data, despite the risk of overfitting when many thousands of teams participate in the same challenge (Roelofs et al., 2019a).
@@ 3 Reconsidering Generalization for GenAI
As we recalled above, the IID assumption in traditional ML evaluations gives a clear conception of the idea of generalization: a model generalizes well if it accurately predicts the true-but-hidden label $y$ for labeled examples $({\bf x},y)$ drawn IID from the same fixed and stationary distribution $D$ from which the model’s training data was drawn. This was a cornerstone allowing the field of ML to progress effectively by narrowing the problem and enabling tractable statistical theory. But if we reflect on broader notions of intelligence, including those first proposed in the seminal paper by Turing (Turing, 1950), it is clear that this narrow-but-useful notion of generalization does not adequately reflect the deeper goals that GenAI is aiming to deliver on.
@@
In order to assess novelty-centric generalization, we need novelty-based evaluations. Informally, the goal of a novelty-based evaluation is to ensure that no evaluation task or example is too closely similar (for some definition of similar and some measure of too close) to any instance previously known to the model or system. We illustrate the distinction between IID-based generalization and novelty-based generalization in Figure 1 visually—we can imagine a small conceptual ring around each training example and ensure that no evaluation instance crosses any of those rings.
In our view, this novelty-centric view of generalization has already been implicitly adopted by many in the field as the true aspirational goal, and influences the design of important benchmarks including the LM Arena (Chiang et al., 2024a) among others—we are simply writing this de facto standard down. We will now examine some of the implications.
@@ 3.2 Leakage and Contamination Are the Biggest Pitfalls
A novelty-centric evaluation rests on assurance that a model has never before been exposed to data that are too close to the evaluation problems or tasks. While this may seem obvious, in practice it can be extremely difficult, as GenAI models like LLMs are often trained on enormous amounts of data and it can be extremely difficult to say for sure what similar data may or may not have been included. Indeed, leakage for GenAI is so important that specific forms of it have been given an additional name: contamination (Magar & Schwartz, 2022; Oren et al., 2023; Sainz et al., 2023; Balloccu et al., 2024). Contamination is said to occur when evaluation datasets and benchmarks appear in training data.
To help give intuition for the breadth of this issue, consider that every major LLM we have tested so far (both open and proprietary) shows extensive detailed knowledge of the contents of standard test datasets from Kaggle. Consider the remarkably strong performance on many static benchmarks by LLMs that do not seem to correlate with strong performance on other tasks (Fourrier et al., 2024; Muennighoff et al., 2023; Zheng et al., 2023). Consider the question: if a model does particularly well on qualification exam normally given to humans, is this because the model has gained strong expertise or because example examinations have appeared in its training data, and how would we be able to disambiguate? Consider the difficulty in teasing apart exactly which data sources are or are not part of an openly shared dataset such as the open and widely used Nectar dataset (Zhu et al., 2024), which includes the description:
@@ 4 Leakage Case Studies
Because leakage and contamination are the most important hurdles to solve for GenAI evaluations, it is useful to study them in depth, beginning with leakage from traditional ML evaluations. Here, we draw on lessons learned surveying more than a decade of Kaggle competitions, in which a broad range of leakage issues have been identified through intense scrutiny of a large community. Experience has shown that the risk of leakage is compounded in open machine learning challenge benchmarks, where teams will exploit (knowingly or unknowingly) anything that gives an advantage on the leaderboard.
@@
Any data that is synthetically generated is highly prone to having artifacts that leak information. Again, in the SETI Breakthrough Listen competition, synthetically-created “ET” signals were injected into real radio telescope signals. Care was taken with normalization to ensure the averages and standard deviations of the injections matched the background signals. The code that created the injected signals used FP16 while the background signals were FP32. This created a minute difference in the mean and standard deviations between positive and negative samples, but enough to differentiate the classes based on this information alone.
Private evaluation data leaking to the public during an open challenge is a risk that needs to be considered. During the LANL Earthquake Prediction challenge (RL et al., 2019), for example, the dataset was described in a research paper, including some summary statistics and a graph. A few teams discovered this and were able to utilize it to their advantage.
@@ 4.1 Reproducibility and Robustness in Conflict
Because of the importance of leakage and the practical difficulty in ensuring that leakage does not impact GenAI evaluations, we argue that it is simplest and safest to adopt a leakage rule of thumb that an evaluation should be considered leaked the moment it has been shared online or sent over the wire. Adopting this rule of thumb significantly improves our ability to trust the results of evaluations and gives them substantially more robustness. However, it also critically weakens the notion of reproducibility. It is the position of this paper that this is a fundamental tension, analogous to the Heisenberg Uncertainty Principle from quantum physics, and that we simply cannot have a published static benchmark that is robust to leakage—no matter the good intentions of the researchers, it is just too hard to avoid contamination and to broadly trust results from such a benchmark.

### v2

@@ Abstract
In this position paper, we observe that empirical evaluation in Generative AI is at a crisis point since traditional ML evaluation and benchmarking strategies are insufficient to meet the needs of evaluating modern GenAI models and systems. There are many reasons for this, including the fact that these models typically have nearly unbounded input and output spaces, typically do not have a well defined ground truth target, and typically exhibit strong feedback loops and prediction dependence based on context of previous model outputs. On top of these critical issues, we argue that the problems of leakage and contamination are in fact the most important and difficult issues to address for GenAI evaluations. Interestingly, the field of AI Competitions has developed effective measures and practices to combat leakage for the purpose of counteracting cheating by bad actors within a competition setting. This makes AI Competitions an especially valuable (but under-utilized) resource. It is now time for the field to view AI Competitions as the gold standard for empirical rigor in GenAI evaluation, and to harness and harvest their results with according value.
@@ 1 Introduction
As Generative AI (GenAI) models such as Large Language Models (LLMs) become ever more important to the field and to the world, it has become clear that performing empirical evaluation of these models and methods is extremely difficult to do in a rigorous and comprehensive way. This difficulty is of course not due to lack of effort or expertise by researchers. Indeed, enormous effort and resources have been poured into creating myriad benchmarks and test cases (Chiang et al., 2024b; Fourrier et al., 2024; Hendrycks et al., 2021; Cobbe et al., 2021; Zellers et al., 2019; Chen et al., 2021b). However, even accounting for these many important efforts and achievements, our position is that the current state of evaluation is insufficient to meet the needs of this moment in GenAI for the field and for the world.
In our view, the root cause of this insufficiency is that the evaluation needs of GenAI models fundamentally break the paradigm of traditional benchmarking that served the field of machine learning (ML) so well during decades of remarkable progress. This breakage goes beyond the familiar difficulty of defining what, exactly, is in the training data for an LLM. In our view, we need a broader conception of generalization for GenAI that moves beyond the idea of generalizing to new independently drawn examples from a stationary distribution, and instead refers to performing well on tasks that are entirely novel from a model’s perspective. This higher bar is rooted in commonsense standards for human intelligence (Chollet, 2019; Dennett, 1991), but has far reaching consequences, most notably that it implies the problems of data leakage and contamination in evaluation are the most pressing concerns.
Together, these factors imply that rigorous and robust evaluation of GenAI models requires a steady source of novel tasks structured to avoid leakage, contamination, and other forms of inadvertent “cheating”. Fortunately, AI Competitions—such as those hosted on platforms like Kaggle and others—act as a solution to GenAI evaluation challenges by providing a continual source of new tasks for evaluation and significant structures to avoid leakage and related issues. We define an AI Competition as a problem or task with an objective evaluation function for ranking solutions or models in which multiple parallel attempts are made during a time-bound period by independent teams.
@@ 1.1 Summarizing Our Position
  • GenAI evaluations should be considered leaked the moment test data has been shared online or sent over the wire to models.
@@
  • Meta-analyses across evaluations should be valued as highly in the field of AI as they are in fields such as medicine.
@@ 1.2 Structure of This Paper
In the remainder of this paper, we will first review the most typical structure and assumptions in traditional ML evaluation and discuss why they are insufficient for GenAI evaluations. We will examine the nature of generalization for GenAI, how this leads to specific concerns around leakage, and additionally show how goals of reproducibility and robustness in evaluation may be fundamentally at odds. We will then show how difficult the problem of leakage is even for traditional ML evaluations with some brief case studies, and look at current GenAI benchmarks that are aiming to overcome leakage and contamination. We finish with an examination of the ways that AI Competitions address these issues, discuss our recommendations, and examine alternate viewpoints. Our goal is to provide convincing support of the view that AI Competitions do indeed provide a gold standard for empirical rigor in evaluating GenAI models, and that the field should place accordingly high value and attention on their results.
Figure 1: IID Evaluations vs. Novelty-Centric Evaluations. In the IID evaluation, left, both training and test data are drawn from the same distribution, resulting in significant overlap in examples in each set. In the novelty-centric version, right, no test example is allowed to be too similar to any given training example. We argue that the latter conceptualization more closely mirrors desired behavior for GenAI evaluations, where generalization is expected to connote the ability to respond well on totally novel inputs.
@@ 2 Background: Revisiting Benchmarking
Traditional ML benchmarking has been founded on the idea of a test-train split, in which an evaluation is structured by training a model from scratch on a given portion of training data and then evaluating that trained model on a holdout set of test data (Mitchell, 1997). This conceptual structure is so fundamental to modern ML practice that it may sometimes be taken for granted. So let us take a moment to examine this basic structure and its implications.
In classical supervised ML, the most common traditional setup is to evaluate a model $f({\bf x})\rightarrow y$, with ${\bf x}\in\Re^{d}$ as feature vectors in some $d$ dimensional feature space and $y\in Y$ as a space of possible labels, such as $\{0,1\}$ for binary classification or $(0,1)$ for regression on probabilities. Labeled examples $({\bf x},y)$ are assumed to have come from some distribution $D$. The training set $D_{train}$ and test set $D_{test}$ are each independently and identically drawn (IID) from $D$, with only the examples in the training set used to fit the model $f({\bf x})$ and only the examples in the test set used to evaluate the model (Mitchell, 1997).
The IID requirement on test-train splits is often taken as a footnote in practice, but in reality it is a cornerstone of the robustness of this setup. The reason for this is that we fundamentally wish our evaluations to be interpretable as statements on the generalization ability of our models: we wish to know how the model will perform on future, previously unseen data. But achieving this is harder than it may sound, because the ML models in question are often of extremely high dimensionality and thus may be prone to overfitting.
@@
A second approach is to use additional data for evaluation. The issue here to be aware of is the classical statistical trap that correlation does not necessarily imply causation, and that trying to assess generalization ability of a model that has no specific mechanism for disambiguating correlation from causal factors may lead to wildly unreliable performance estimates. It is this issue that the IID assumption addresses. When we know that all test data is drawn IID from the same distribution as the training data, we know that all correlations that held at training time will reappear with the same characteristics at test time, and thus we can take performance on holdout test data as a reasonable estimate of generalization ability. The IID assumption has, in many ways, enabled modern ML research to advance as a field, because it forms the theoretical underpinning of all evaluations. And indeed, it is a truism that moving ML models from research to deployed production is difficult precisely because of the fact that the IID assumption often does not hold in practice (Chen et al., 2021a).
@@ 2.1 The Rise of Reproducible Benchmarks
One statistical shortcut that immediately became standard practice was instead of drawing a new training set and test set from $D$ for every new evaluation, researchers would make one draw of each and use them as a canonical train / test set. The primary benefit of this approach (besides convenience) was that these paired test-train splits could now be used as reproducible benchmarks. All future researchers could replicate the exact problem setup, giving a new untrained model the same training data for training and the same test data for evaluation, allowing for full apples-to-apples comparisons. This approach was wildly successful, with canonical benchmarks such as MNIST (LeCun & Cortes, 2010) and ImageNet (Deng et al., 2009) responsible for driving incredibly rapid progress in computer vision, for instance, and benchmarks such as those of Rajpurkar et al. (2016); Marcus et al. (1993); Diemert Eustache, Betlei Artem et al. (2018) moving forward email spam classification, natural language processing, and many other domains. Websites such as the UCI Machine Learning Repository (Kelly et al., 2025) and OpenML (Vanschoren et al., 2014) among many others have been remarkably valuable to the field for this reason.
@@ 2.2 Surprisingly, Overfitting Was Not the Main Issue
Given that one of the fundamental concerns in evaluating models was ensuring that we could properly address generalization and avoid overfitting, it is reasonable to ask whether widespread reuse of the same standard benchmark datasets in thousands of papers might not lead to overfitting. As authors, we were deeply surprised by the work of Roelofs et al. (2019b) showing that in practice, this appears to actually not have been a problem. Recht et al. (2019) shows that the rank ordering of ImageNet models when evaluated on brand new data was remarkably consistent with the rank ordering of those models on the benchmark data, despite wide reuse. And in follow-up work, Roelofs et al. (2019a) similarly show that evaluation on public leaderboard data on Kaggle competitions was a remarkably good indicator of rank ordering on private holdout data, despite the risk of overfitting when many thousands of teams participate in the same challenge.
@@ 3 Reconsidering Generalization for GenAI
As we recalled above, the IID assumption in traditional ML evaluations gives a clear conception of the idea of generalization: a model generalizes well if it accurately predicts the true-but-hidden label $y$ for labeled examples $({\bf x},y)$ drawn IID from the same fixed and stationary distribution $D$ from which the model’s training data was drawn. This was a cornerstone allowing the field of ML to progress effectively by narrowing the problem and enabling tractable statistical theory. But if we reflect on broader notions of intelligence, including those first proposed in the seminal paper by Turing (1950), it is clear that this narrow-but-useful notion of generalization does not adequately reflect the deeper goals that GenAI is aiming to deliver on.
@@
In order to assess novelty-centric generalization, we need novelty-based evaluations. Informally, the goal of a novelty-based evaluation is to ensure that no evaluation task or example is too closely similar (for some definition of similar and some measure of too close) to any instance previously known to the model or system. We illustrate the distinction between IID-based generalization and novelty-based generalization in Figure 1 visually. We can imagine a small conceptual ring around each training example and ensure that no evaluation instance crosses any of those rings.
In our view, this novelty-centric view of generalization has already been implicitly adopted by many in the field as the true aspirational goal, and influences the design of important benchmarks including the LM Arena (Chiang et al., 2024a) among others; we are simply writing this de facto standard down. We will now examine some of the implications.
@@ 3.2 Leakage and Contamination Are the Biggest Pitfalls
A novelty-centric evaluation rests on assurance that a model has never before been exposed to data that is too close to the evaluation problems or tasks. While this may seem obvious, in practice it can be extremely difficult, as GenAI models like LLMs are often trained on enormous amounts of data and it can be extremely difficult to say for sure what similar data may or may not have been included. Indeed, leakage for GenAI is so important that specific forms of it have been given an additional name: contamination (Magar & Schwartz, 2022; Oren et al., 2023; Sainz et al., 2023; Balloccu et al., 2024). Contamination is said to occur when evaluation datasets and benchmarks appear in training data.
To help give intuition for the breadth of this issue, consider that every major LLM we have tested so far (both open and proprietary) shows extensive detailed knowledge of the contents of standard test datasets from Kaggle. Consider the remarkably strong performance on many static benchmarks by LLMs that do not seem to correlate with strong performance on other tasks (Fourrier et al., 2024; Muennighoff et al., 2023; Zheng et al., 2023). Consider the question: if a model does particularly well on qualification exam normally given to humans, is this because the model has gained strong expertise or because example examinations have appeared in its training data, and how would we be able to disambiguate? Consider the difficulty in teasing apart exactly which data sources are or are not part of an openly shared dataset such as the widely used Nectar dataset (Zhu et al., 2024), which includes the description:
@@ 4 Leakage Case Studies
Because leakage and contamination are the most important hurdles to solve for GenAI evaluations, it is useful to study them in depth, beginning with leakage from traditional ML evaluations. Here, we draw on lessons learned surveying more than a decade of Kaggle competitions, in which a broad range of leakage issues have been identified through intense scrutiny of a large community. Experience has shown that the risk of leakage is compounded in open ML challenge benchmarks, where teams will exploit (knowingly or unknowingly) anything that gives an advantage on the leaderboard.
@@
Any data that is synthetically generated is highly prone to having artifacts that leak information. Again, in the SETI Breakthrough Listen competition, synthetically-created “ET” signals were injected into real radio telescope signals. Care was taken with normalization to ensure the averages and standard deviations of the injections matched the background signals. However, the code that created the injected signals used FP16 while the background signals were FP32. This created a minute difference in the mean and standard deviations between positive and negative samples, but enough to differentiate the classes based on this information alone.
Private evaluation data leaking to the public during an open challenge is a risk that needs to be considered. During the LANL Earthquake Prediction challenge (RL et al., 2019), for example, the test dataset was described in a research paper, including some summary statistics and a graph. A few teams discovered this and were able to utilize it to their advantage.
@@ 4.1 Reproducibility and Robustness in Conflict
Because of the importance of leakage and the practical difficulty in ensuring that leakage does not impact GenAI evaluations, we argue that it is simplest and safest to adopt a leakage rule of thumb that an evaluation should be considered leaked the moment it has been shared online or sent over the wire. Adopting this rule of thumb significantly improves our ability to trust the results of evaluations and gives them substantially more robustness. However, it also critically weakens the notion of reproducibility. It is the position of this paper that this is a fundamental tension, analogous to the Heisenberg Uncertainty Principle from quantum physics, and that we simply cannot have a published static benchmark that is robust to leakage. No matter the good intentions of the researchers, it is just too hard to avoid contamination and to broadly trust results from such a benchmark.